OpenStudy (christos):

Can someone explain to me why sec10* / csc10* = 1? ==> sec10* = csc10*

OpenStudy (anonymous):

\[\frac{a}{b}=1\iff a=b\]

Directrix (directrix):

sec(10*) = csc(10*) --> This is a false statement. sec(10*) = csc(80*) --> They are cofunctions but not reciprocal functions

OpenStudy (anonymous):

good point

OpenStudy (christos):

But if sec10* / csc10* = 1 is True Then why not sec10* = csc10*

Directrix (directrix):

Why do you think this is true: sec10* / csc10* = 1 ?

OpenStudy (christos):

Are you telling me that it is not?

Directrix (directrix):

I am asking you why you think it is true. ------------------ sec(10) / csc(10) = tan(10)

OpenStudy (christos):

Alright then find the exact value of: - sec10* / csc 80*

OpenStudy (anonymous):

secant and cosecant are cofunctions

OpenStudy (christos):

What do you mean by that satellite? :)

OpenStudy (anonymous):

lets look at a triangle 

OpenStudy (anonymous):

you can see from the triangle that \(\sec(10)=\frac{h}{b}\) and also that \(\csc(80)=\frac{h}{b}\) they are the same

OpenStudy (anonymous):

that is because \(10+80=90\) or \(80=90-10\) which ever you prefer so you can see from the picture that in general \[\sec(\theta)=\csc(90-\theta)\]

OpenStudy (anonymous):

same works for cosine and sine and cotangent and tangent that is why they are called "co - functions"

OpenStudy (anonymous):

so since \(\sec(10)=\csc(80)\) you also have \[\frac{\sec(10)}{\csc(80)}=1\]

OpenStudy (anonymous):

in all fairness this is not what you wrote in the question you had \(\frac{\sec(10)}{\csc(10)}=1\)which is false

OpenStudy (christos):

I was marely trying yo understand the concept of it

OpenStudy (christos):

to*

OpenStudy (christos):

@satellite73 Can you please solve for me - sec10*/csc80* ?

Directrix (directrix):

-sec(10) divided by csc(80) = - 1